# Contextual Bandit with Herding Effects: Algorithms and Recommendation Applications

### v1

@@ 2 Related Work
Contextual linear bandits serve as a fundamental sequential decision making framework for information retrieval applications advertising, recommendation, etc [5]. A number of variants of contextual linear bandits were proposed to capture important factors of information retrieval applications [24, 25, 26, 34]. Conversational contextual bandit tailors the contextual linear bandit to capture conversational feedbacks in recommendation applications [32]. The contextual budgeting bandit extend the contextual linear bandit to the multi-agent for the purpose of studying the budget allocation in online advertising [6]. The key difference to the above work is that our model captures the well-known herding effects in feedback. The new technical challenge is that this herding effects leads to confounded feedback with spurious correlation, which may result in linear regret.
@@ 4.1 Algorithm Design
$p\left(\bm{\Psi}\mid\mathcal{H}_{t}\right)=$ $\frac{p(\bm{\Psi})}{C}\times\left[\prod_{\tau=1}^{t-1}\prod_{a\in\mathcal{A}_{\tau}}\left[f\left(\eta_{\tau,a},\sigma_{a}\right)\right]^{\mathbbm{1}_{\left\{A_{\tau}=a\right\}}}\right],$ (4)
@@ 4.2 Regret Analysis
$R_{T}(\mathcal{D})$ $\leq O\left(\frac{1}{1-\alpha}\int\sum^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})d\bm{\Psi}\right.+\left.\frac{1}{1-\alpha}d\sqrt{T}\ln T\right).$
@@
The above theorem shows that the algorithm can effectively learn from historical decisions and gradually approach the performance of the best possible decision. Due to page limitations, more details on the proofs of the above two theorems can be found in the appendix.
@@ 5.1 Experiment Settings
Datasets. Our approach to constructing the simulated datasets is aligned with established practices in the field, similar to those employed in related studies[30, 13]. In our empirical evaluation using real-world data, we employ datasets sourced from four distinct platforms: Amazon Music [^1] http://jmcauley.ucsd.edu/data/amazon/index\_2014.html, MovieLens [^2] https://grouplens.org/datasets/movielens/, Yelp [^3] https://www.yelp.com/dataset, and Google Maps [^4] http://jmcauley.ucsd.edu/data/googlelocal/. Through the utilization of these varied datasets spanning multiple domains, our objective is to rigorously assess and understand the real-world applicability and performance of the proposed algorithm. Based on the approach similar to [30, 33], we perform data preprocessing and assess the accuracy of algorithm recommendations through regret values. Due to page limitations, more details on the data preprocessing for datasets can be found in the appendix.
@@ 5.3 Real-world applications
Results in MovieLens. Figure 4 shows the regret $\hat{R}_{t}$ produced by each algorithm in different dimensions and different noise variances. It can be observed that the TS-Conf algorithm always has the lowest regret value across varying dimensions and noise levels. Consistent with the experimental observations on synthetic data, the regret values for both LinUCB and TS (both of them ignore the biased influence in user feedback) exhibit significantly higher regret than the TS-Conf algorithm and will increase linearly with $t$. Similarly, LinUCBConf also has a greater regret compared with TS-Conf, and its regret values converge at a slower rate than our algorithm.
@@
Results in Yelp. Figure 5 shows the comparison results of the four algorithms on the Yelp dataset. It is evident that the TS-Conf algorithm consistently has the lowest regret value across different dimensions and noise levels. Unlike the LinUCB and TS algorithms, which consistently exhibit linear regret growth, the regret of TS-Conf gradually converges over time. Furthermore, its convergence speed is greater than that of LinUCBConf. Similar results for Amazon Music and Google Maps datasets are presented in the appendix due to page limitations.
@@ 6 Conclusion
This paper presents a contextual bandit framework to address the herding effects problem in recommendation applications. In this framework, we assume that the user feedback on the action is biased and influenced by user preferences and the historical ratings of this action. We design a TS-Conf algorithm that leverages a posterior sampling technique to effectively balance the trade-off between exploration and exploitation in this framework. Our theoretical analysis established a sublinear regret bound for the TS-Conf algorithm, demonstrating its efficiency. Extensive experiments on synthetic and real-world datasets, including those from Amazon, MovieLens, Yelp, and Google, show that TS-Conf consistently outperforms three benchmark algorithms, confirming its robustness and superior performance in handling herding effect-induced biases.

### v2

@@ 2 Related Work
Contextual linear bandits serve as a fundamental sequential decision making framework for information retrieval applications advertising, recommendation, etc [5]. A number of variants of contextual linear bandits were proposed to capture important factors of information retrieval applications [25, 26, 27, 35]. Conversational contextual bandit tailors the contextual linear bandit to capture conversational feedbacks in recommendation applications [33]. The contextual budgeting bandit extend the contextual linear bandit to the multi-agent for the purpose of studying the budget allocation in online advertising [6]. The key difference to the above work is that our model captures the well-known herding effects in feedback. The new technical challenge is that this herding effects leads to confounded feedback with spurious correlation.
@@ 4.1 Algorithm Design
$p\left(\bm{\Psi}\mid\mathcal{H}_{t}\right)=$ $\frac{p(\bm{\Psi})}{C}\times\left[\prod\nolimits_{\tau=1}^{t-1}\prod\nolimits_{a\in\mathcal{A}_{\tau}}\left[f\left(\eta_{\tau,a},\sigma_{a}\right)\right]^{\mathbbm{1}_{\left\{A_{\tau}=a\right\}}}\right],$ (4)
@@ 4.2 Regret Analysis
$R_{T}(\mathcal{D})$ $\leq O\left(\frac{1}{1-\alpha}\int\sum\nolimits^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})d\bm{\Psi}\right.+\left.\frac{1}{1-\alpha}d\sqrt{T}\ln T\right).$
@@
The above theorem shows that the algorithm can effectively learn from historical decisions and gradually approach the performance of the best possible decision. Due to page limitations, more details on the proofs of the above two theorems can be found in our technical report [11].
@@ 5.1 Experiment Settings
Datasets. Our approach to constructing the simulated datasets is aligned with established practices in the field, similar to those employed in related studies[31, 14]. In our empirical evaluation using real-world data, we employ datasets sourced from four distinct platforms: Amazon Music [^1] http://jmcauley.ucsd.edu/data/amazon/index\_2014.html, MovieLens [^2] https://grouplens.org/datasets/movielens/, Yelp [^3] https://www.yelp.com/dataset, and Google Maps [^4] http://jmcauley.ucsd.edu/data/googlelocal/. Through the utilization of these varied datasets spanning multiple domains, our objective is to rigorously assess and understand the real-world applicability and performance of the proposed algorithm. Based on the approach similar to [31, 34], we perform data preprocessing and assess the accuracy of algorithm recommendations through regret values. Due to page limitations, more details on the data preprocessing can be found in our technical report [11].
@@ 5.3 Real-world applications
Results in MovieLens. Figure 4 shows the regret $\hat{R}_{t}$ produced by each algorithm in different dimensions and different noise variances. It can be observed that the TS-Conf algorithm always has the lowest regret value across varying dimensions and noise levels. Consistent with the experimental observations on synthetic data, the regret values for both LinUCB and TS exhibit significantly higher regret than the TS-Conf algorithm and will increase linearly with $t$. Similarly, LinUCBConf also has a greater regret compared with TS-Conf, and its regret values converge at a slower rate than our algorithm.
@@
Results in Yelp. Figure 5 shows the comparison results of the four algorithms on the Yelp dataset. It is evident that the TS-Conf algorithm consistently has the lowest regret value across different dimensions and noise levels. Unlike the LinUCB and TS algorithms, which consistently exhibit linear regret growth, the regret of TS-Conf gradually converges over time. Furthermore, its convergence speed is greater than that of LinUCBConf. Similar results for Amazon Music and Google Maps datasets are presented in our technical report[11].
@@ 6 Conclusion
This paper presents a contextual bandit framework to address the herding effects problem in recommendation applications. In this framework, we assume that the user feedback on the action is biased and influenced by user preferences and the historical ratings of this action. We design a TS-Conf algorithm that leverages a posterior sampling technique to effectively balance the trade-off between exploration and exploitation. Our theoretical analysis established a sublinear regret bound for the TS-Conf algorithm, demonstrating its efficiency. Extensive experiments on synthetic and real-world datasets show that TS-Conf consistently outperforms three benchmark algorithms, confirming its robustness and superior performance in handling herding effect-induced biases.